OpenStudy (anonymous):

The Andersons traveled 90 miles in 75 minutes. At this rate, how far would they travel in 2 hours?

OpenStudy (anonymous):

90/75 = x/120 x=90(120)/75=144 miles

OpenStudy (anonymous):

Thank you!